the motion of a particle is given by x(t)=(25cm)cos(12t) where t is in s

The motion of a particle is given by x(t)=(25cm)cos(12t), where t is in s.

What is the first time the kinetic energy is twice the potential energy?
Answer #1

when x=A*sqrt(3)/3 or


1/2kA^2 = 1/2kX^2+2*1/2kX^2=3/2kx^2
x^2=A^2/3 .....
x=A/sqrt(3)=A*sqrt(3)/3

cos(12t)=sqrt(3)/3

t=0.0796 s